Find the exact length of the side of a square whose diagonal is 3 feet.
Answer by checkley77(12569)

X^2+X^2=3^2
2X^2=9
X^2=9/2
X=SQRT4.5
X=2.12132 FT. IS THE LENGTH OF THE SIDE OF THE SQUARE.